# Evaluating LLMs on microservice-based applications: how complex is your specification?

### v1

@@ Abstract
In this paper we evaluate how far LLMs have advanced in generating code for real-world problems. Specifically, we explore code synthesis for microservice-based applications, a widely used architecture pattern. We define a standard template for specifying these applications, and we propose a metric for judging the difficulty level of a specification. The higher the score, the more difficult it is to generate code for the specification. We develop a framework to automate the process of testing the LLM-synthesized code for a microservice using unit tests. Our experimental results show that strong LLMs (like GPT-3o-mini) do fairly well on medium difficulty specifications but do very poorly on those of higher difficulty levels. This is due to more intricate business logic, a greater use of external services, database integration and inclusion of non-functional capabilities such as authentication. We analyzed the errors in LLM-synthesized code and report on the key challenges LLMs face in generating code for these specifications thereby suggesting future research directions to improve code synthesis for real-world problems.
@@ 1 Introduction
The rapid maturation of Large Language Models (LLMs) for code generation has widely influenced software development, with LLMs now routinely used for software engineering tasks including code generation. Benchmarks have become the de facto standard for measuring an LLMs’ performance. Early benchmarks, such as HumanEval[2] and MBPP[1] determine how well LLMs perform in generating a single function of increasing algorithmic complexity. State-of-the-art models (including reasoning models such as GPT-o1 and GPT-o3 and DeepSeek-R1) show impressive results against these benchmarks, with results approaching 100% for HumanEval [8].
However, these benchmarks shed little light on how well these models can perform on real-world problems, where the complexity is of a different nature. In real software development, complexity expresses itself not only in the computation (algorithm) of a particular component, but in the integration of different software components and in the distributed nature of modern systems; i.e., Cloud Native architectures [21]. Newer benchmarks, such as SWE-Bench [10] have been developed to address these more realistic scenarios by asking LLMs to fix a specific bugs in real GIT repos [33, 17, 31, 3] or to synthesize code to add a new feature/function to a repo [13, 32, 15].
The main challenge that GIT repo benchmarks poise to LLM code generation is the size of the context. Popular GIT repos often have hundreds of thousands LOC [10] which is too large to fit into LLM context windows. Hence newer techniques, including agentic solutions, often include information retrieval techniques and focus on how to extract the right context to feed the LLM [30, 28, 29].
There is an orthogonal problem in generating software for real-world problems that has received less attention; namely, how to specify a complete solution to be generated. Real-world software is complex and requires far more descriptive text than required for fixing a GIT repo issue. How well do LLMs perform in this scenario?
To make the problem concrete, we focus on LLM code synthesis for a core design pattern of modern systems, Microservice-Based Applications (MSBAs). These apps have the following characteristics:
@@
The GitHub repo problem focuses on code generation where the code base is complex and cannot fit into the context window of the LLM. The MSBA problem, in contrast, focusses on code generation from scratch, where the specification can easily fit into the context window of the LLM, but the challenge is in the complexity of the spec; i.e., the intricacies of the software being developed, not code size. For instance, SWE-Bench issues provided to an LLM to fix are 195 words on average [^1] However, the max length for SWE-Bench issues is 4477 words. [10]. Our MSBA examples have descriptions of 1399 and 1905 words. While these are not full-fledged production-level apps, they serve as a good testbed for our experiments. In Section 3.2 we discuss metrics for measuring the complexity of MSBAs. As far as we know, there are no existing benchmarks for MSBAs, and little if any work has examined using LLMs to generate such systems.
The purpose of this paper is to provide a preliminary study on how well LLMs perform in this important domain, and to encourage more research to address the challenges we find.
@@ 2 Challenges in generating MSBAs
Some of challenges in synthesizing MSBAs from LLMs are:
Integration: includes reconciling type and naming mismatches between different frameworks and libraries.
@@
Specification: There is no standard for specifying a MSBA to an LLM. Unlike simple functions, a docstring and a signature do not suffice.
On the other hand, MSBAs using REST services often follow a well-established template (see Appendix C for a small example) giving hope that LLMs can exploit this template to generate correct code.
@@
  • How should one specify a MSBA to facilitate LLM code generation and how can one measure the complexity of a specification? This is discussed in Section 3.
@@
  • How well do LLMs perform in generating MSBAs? How does the complexity of the specification affect LLM performance? Our experimental results are given in Section 5.
3 Specifying a MSBA
To describe an MS that is part of a MSBA you need to identify:
@@ 3 Specifying a MSBA
  • The requests (HTTP verbs) that can be issued on these resources. The basic verbs are the CRUD-like requests POST, GET, PUT and DELETE. Each operation will specify additional details. For instance, a GET request must specify its payload (parameters), allowable query strings, and details specific to HTTP and RESTful APIs, such as media types accepted and status codes to be returned. The spec may also limit whom is allowed to issue requests.
@@
{"Name": "Cardholders microservice","Description": "Responsible for creating, retrieving, updating, and deleting     cardholder accounts. This service also provides the ability to retrieve     how much money in fines is owed by a specific cardholder.”"Resources": "/cardholders, /cardholders/{id},/cardholders/fines/{id}. An id is     of type string. Each cardholder (/cardholders/{id}) consists of a JSON     document containing the field names ‘name’, ‘email’, and ‘id’. The values     of each of these is a JSON string. The resource /cardholders/fines/{id}     consists of a JSON document with a field ‘fineAmount’, whose value is     of type float, and a field ‘id’, of type string.","REST requests": "GET and POST request on /cardholders. GET returns a     JSON array of all records in the resource. The GET request may specify a     query string of the form ‘field=value’. In this case, only those cardholder     records satisfying that constraint should be returned. The POST request     provides a JSON document containing the name and email and returns the     id for the newly created record.     GET,PUT and DELETE requests on /cardholders/{id}. The GET request     returns a JSON record of the requested resource.  The PUT request provides     a JSON payload of the updated JSON resource’s fields and returns the id of     the updated resource. The DELETE request returns the status of the request     with no content.     GET request on cardholders/fines/{id}.  The request returns the JSON     document representing this resource.","Additional details”: "The fine amount owed is computed as follows: at any given     day, the fine per overdue book is equal to number-of-days-overdue *     FINE_PER_DAY, where FINE_PER_DAY equals \$.50. The cardholder owes     the sum of all fines for all books he borrowed. To compute this amount,     this service needs to invoke the Borrows MS to retrieve information on books     borrowed by this cardholder.","Deployment": "The Cardholders MS runs in a Docker container and listens on     external port 5001. Another MS can send it requests using the URL     http://cardholders:5001/cardholders."}
@@
In Figure 1 we illustrate the specification of a Cardholders MS, part of a simple public library application we present in Section 5.3. “Description” provides the purpose of the MS. “Resources” lists the REST resources that the service is responsible for. “REST requests” defines the type of REST requests supported for each resource, and how resources are to be represented. “Additional details” provides functional requirements for the service. “Deployment” states deployment details; for instance, that the micro-service will run in a Docker container, the port the MS listens on and the URL used to send this MS requests. An alternative to this format would be to use OpenAPI (Swagger) specifications [18] but we would need to add information necessary to implement the server implementation. The format we provide is more compact.
All prompts we use are given in Appendix A, including the LLM Code Generation Guidelines. These guidelines instruct the LLM to build the MSes using REST services that execute Python code, and that store data in a MongoDB. It also defines which HTTP status codes should be returned in case of normal or exceptional execution.
@@ 3.1 Creating MSBA specifications
To evaluate LLMs on MSBAs, we created 8 MSes across 2 MSBAs. We created our own MSBAS for two reasons. First we want to make sure that the LLMs have not been contaminated - have not seen these MSes before. Second, we want to compare LLM’s code generation capabilities on MSBAs of increasing specification complexity, but still small enough to focus on core MSBA challenges (Section 2) and not on extraneous issues. We built the MSBA specifications using the following process:
@@
    When testing and finding an error, we checked the cause of the error. (a) If it was a mistake in the GT implementation, we fixed the GT implementation. (b) If the test case was not faithful to the spec, we corrected the test case. (c) If the spec was ambiguous or incorrect, we fixed the spec.
@@
The specifications and list of unit tests for each MSBA can be found online222 https://github.com/danny-m-yellin/LLMs4MSBA-artifacts.
@@ 3.2 The complexity of an MSBA specification
We measure the complexity of a a MSBA in several ways:
@@
The specification size indicates the number of details the LLM must contend with. The number of dependencies and packages indicates the amount of external knowledge required by the LLM. These metrics are not exact; its possible to have a large specification that is simple to implement and a small specification that is difficult to implement. Similarly not all external dependencies are equal; some express more intricate interfaces than others. Nonetheless, these metrics give a rough estimate on the complexity of an MSBA spec and are easy to compute. The LLM-as-a-Judge serves as a validation for the other metrics. Furthermore, the LLM explains why it gives the score, which provides further insight into the complexity of the spec.
@@ 4 Testing MSBAs
The result of a request to a MS is sensitive to history. For instance, the outcome of a GET request is dependent on the history of the preceding POST, PUT and DELETE requests. One needs to test multiple scenarios for a GET to determine whether it functions correctly. Furthermore, the unit tests must also check that the appropriate exceptions are raised when warranted; e.g., when a GET request is issued on an item previously deleted, the status code 404 must be returned whereas if the request provides the wrong media type a status code 415 is returned. Developing adequate unit tests can be time consuming and tedious. To help simplify and make testing of REST APIs more consistent, we developed a REST API testing framework, that consists of 4 modules.
@@
The main module is (i) Unit Test, which issues the unit tests. It relies on the (ii) HTTP Service Controller to format and invoke the requests according to the HTTP protocol. We divide unit tests into two sets: common unit tests and MS-specific unit tests. Common tests are those where the result depends on the pattern of the requests, without concern for their specific input and output values. For instance, if we issue requests POST $X_{1}$, POST $X_{2}$, POST $X_{3}$, DELETE $X_{2}$ and finally GET, we expect the return value to be the JSON array containing $X_{1}$ and $X_{3}$ and nothing else. This is true independent of the actual values of $X_{i}$. We rely on this insight to parameterize unit tests. Parameter values, expected returned values, and any other MS specific values needed to test the MS are defined in the (iii) Database module. The Common Unit Test module fetches the appropriate values from the DB when testing a MS; for the example above, each MS will provide different values for $X_{1}$, $X_{2}$ and $X_{3}$.
@@
The common unit tests may still require some customization, and not every custom test applies to all MSes. However, the framework greatly simplifies building tests for a new MS. Likewise, we created a reusable parameterized (iv) assertion library that checks return values satisfy assertions. For example, the assertion assert-fields-contains-value(response: requests.Response,f:str,v) asserts that the response is a JSON array, each element in the array is a struct with field f, the field f is itself a JSON array, and f contains the value v.
@@
In addition to common unit tests, we need to test MS-specific functionality. For example, we want to check that the Cardholders MS (Figure 1) correctly computes the fine for overdue books. MS-specific tests need to be customized per MS.
@@
Say that an MSBA is made up of 4 MSes: $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. Let $M^{GT}_{i}$ be the GT implementation of $M_{i}$ and $M^{LLM}_{i}$ be the implementation of $M_{i}$ generated by the LLM. To test $M^{LLM}_{2}$ we run the MSBA using the MS implementations $M^{GT}_{1}$, $M^{LLM}_{2}$, $M^{GT}_{3}$, and $M^{GT}_{4}$ and then test the application. In this way we isolate the test failures due only to $M^{LLM}_{2}$.
@@ 5.1 Methodology
Given an MSBA specification, we ask the LLM to generate the MSs of that MSBA. Our implementation is a workflow (WF) built on top of LangGraph [11], as illustrated in Figure 2. After (1) initializing the WF, it (2) prompts the LLM with the MSBA description including code generation guidelines, asking it to generate one MS $M$. The prompt is given in Appendix A. It then (3) creates a requirements.txt file for the generated code. (4) Next it sets up the Docker containers to execute the LLM generated code $M$ and the GT implementations for all other MSes and then (5) executes a Docker Compose script to execute the unit tests.
Figure 2: Workflow for generating and testing an MSBA. The blue arrows indicate control flow. The dotted black arrows indicate writing and reading data. Blue rectangular boxes are LLM interactions.
(6) Once the tests complete, or the execution is interrupted by run-time errors, the containers are shutdown and (7) the error messages are examined. If there are no runtime and test errors, the LLM synthesized code is successful and the workflow (11) records the results. (8) Otherwise it checks if it has reached the maximum number of code-generation trials. If not, (9) the WF asks the LLM to reflect [23] on the errors and produce a plan to fix the generated code to prevent the top 5 errors. This prompt is given in Appendix A. Because of context window limitations of LLMs, and because the runtime may produce a very long trace of each error, we use a script to remove parts of the error messages that are not useful to debug the issue. For some LLMs, we need to further limit the size of the error messages to fit into the context window. (10) The original code together with the result of the reflection is then provided to the LLM, which is instructed to generate code that fixes the problems. This prompt is given in Appendix /refappendix-prompts. The workflow then loops back to step (3). The entire WF repeats for each MS.
In summary, our workflow automates the entire process from generation to testing to feedback by invoking multiple agents. It uses LangGraph to orchestrate the flow, runs the generated code safely in Docker containers, and uses reflection to improve results.

### v2

# LLM Agents for Generating Microservice-based Applications: How Complex is Your Specification?

@@ Abstract
In this paper we evaluate the capabilities of LLM Agents in generating code for real-world problems. Specifically, we explore code synthesis for microservice-based applications, a widely used architectural pattern for building applications. We define a standard template for specifying these applications, and we propose a metric for scoring the difficulty of a specification. The higher the score, the more difficult it is to generate code for the specification. Our experimental results show that agents using strong LLMs (like GPT-3o-mini) do fairly well on medium difficulty specifications but do poorly on those of higher difficulty levels. This is due to more intricate business logic, a greater use of external services, database integration and inclusion of non-functional capabilities such as authentication. We analyzed the errors in LLM-synthesized code and report on the key challenges LLM Agents face in generating code for these specifications. Finally, we show that using a fine-grained approach to code generation improves the correctness of the generated code.
@@ 1 Introduction
The rapid maturation of Large Language Models (LLMs) for code generation has widely influenced software development, with LLM Agents now routinely used for software engineering tasks including code generation. Benchmarks have become the de facto standard for measuring an LLM Agent’s performance. Early benchmarks, such as HumanEval[7] and MBPP[3] determine how well LLM Agents perform in generating a single function of increasing algorithmic complexity. State-of-the-art models (including reasoning models such as GPT-o1, GPT-o3 and DeepSeek-R1) show impressive results against these benchmarks, with results approaching 100% for HumanEval [11].
However, these benchmarks shed little light on how well these models can perform on real-world problems where the complexity is of a different nature. In real software development, complexity expresses itself not only in the computation (algorithm) of a particular component, but in the integration of different software components and in the distributed nature of modern systems; i.e., Cloud Native architectures [25]. Newer benchmarks, such as SWE-Bench [13] have been developed to address realistic scenarios by asking LLMs to fix a specific bugs in real GIT repos [37, 21, 35, 2] or to synthesize code to add a new feature/function to a repo [17, 36, 19].
The main challenge that GIT repo benchmarks poise to LLM code generation is the size of the context. Popular GIT repos often have hundreds of thousands LOC [13] which is too large to fit into LLM context windows. (The average length of 500 real-world repositories is 1.1 million tokens[16].) Hence newer techniques, including agentic solutions, often include information retrieval techniques and focus on how to extract the right context to feed the LLM [34, 32, 33].
There is an orthogonal problem in generating software for real-world problems that has received less attention; namely, how to specify a complete solution to be generated. Real-world software is complex and requires far more descriptive text than required for fixing a GIT repo issue. How well do LLM Agents perform in this scenario?
To make the problem concrete, we focus on using LLM agents to synthesis code for a core design pattern of modern systems, Microservice-Based Applications (MSBAs). These apps have the following characteristics:
@@
The GitHub repo problem focuses on code generation where the code base is complex and cannot fit into the context window of the LLM. The MSBA problem, in contrast, focusses on code generation from scratch, where the specification can easily fit into the context window of the LLM, but the challenge is in the complexity of the spec; i.e., the intricacies of the software being developed, not code size. For instance, SWE-Bench issues provided to an LLM to fix are 195 words on average [^1] The max length for SWE-Bench issues is 4477 words. [13]. Our MSBA examples have descriptions of 1399 and 1905 words. While these are not full-fledged production-level apps, they serve as a good testbed for our experiments. As far as we know little if any research has investigated using agentic software to generate MSBAs.
The purpose of this paper is to provide insight on how well LLM Agents perform in this domain, analyze the types of errors they encounter, and show that the use of fine-grained code generation enhances performance. We also suggest a metric for rating the complexity of a specification.
@@ 2 Challenges in generating MSBAs
Some of challenges in synthesizing MSBAs from LLM Agents are:
Integration: reconciling type and naming mismatches between different frameworks and libraries.
@@
Specification: There is no standard for specifying an MSBA to an LLM. Unlike simple functions, a docstring and a signature do not suffice.
On the other hand, MSBAs using REST services often follow a well-established pattern (see Appendix C for a small example) giving hope that LLM Agents can exploit this pattern to generate correct code.
@@
  • How should one specify an MSBA to facilitate LLM code generation and how can one measure the complexity of a specification? This is discussed in Section 3.
@@
  • How well do LLM Agents perform in generating MSBAs? How does the complexity of the specification affect LLM performance? Our experimental results are given in Section 5.
  • Can a fine-grained approach to code generation can improve the accuracy of the generated code? We show our results in Section 6.
Sections 7 discusses related work and Section 8 presents our conclusion.
3 Specifying an MSBA
To describe a microservice (MS) that is part of an MSBA you need to identify:
@@ 3 Specifying an MSBA
  • The requests (HTTP verbs) that can be issued on these resources. The basic verbs are the CRUD requests POST, GET, PUT and DELETE. Each request will specify additional details. For instance, a GET request must specify its payload (parameters), allowable query strings, and details specific to HTTP and RESTful APIs, such as media types accepted and status codes to be returned. The spec may also limit whom is allowed to issue requests.
@@
{"Name": "Cardholders microservice","Description": "Responsible for creating, retrieving, updating, and deleting cardholder accounts. This service also provides the ability to retrieve how much money in fines is owed by a specific cardholder.”"Resources": "/cardholders, /cardholders/{id},/cardholders/fines/{id}. An id is of type string. Each cardholder (/cardholders/{id}) consists of a JSON document containing the field names ‘name’, ‘email’, and ‘id’. The values of each of these is a JSON string. The resource /cardholders/fines/{id} consists of a JSON document with a field ‘fineAmount’, whose value is of type float, and a field ‘id’, of type string.","REST requests": "GET and POST request on /cardholders. GET returns a JSON array of all records in the resource. The GET request may specify a query string of the form ‘field=value’. In this case, only those cardholder records satisfying that constraint should be returned. The POST request provides a JSON document containing the name and email and returns the id for the newly created record. GET,PUT and DELETE requests on /cardholders/{id}. The GET request returns a JSON record of the requested resource. The PUT request provides a JSON payload of the updated JSON resource’s fields and returns the id of the updated resource. The DELETE request returns the status of the request with no content. GET request on cardholders/fines/{id}. The request returns the JSON document representing this resource.","Additional details”: "The fine amount owed is computed as follows: the fine per overdue book is equal to number-of-days-overdue * FINE_PER_DAY, where FINE_PER_DAY equals \$.50. The cardholder owes the sum of all fines for all books he borrowed. To compute this amount, this service needs to invoke the Borrows MS to retrieve information on books borrowed by this cardholder.","Deployment": "The Cardholders MS runs in a Docker container and listens on external port 5001. Another MS can send it requests using the URL http://cardholders:5001/cardholders."}
@@
In Figure 1 we provide the specification of a Cardholders MS, part of a public library application we present in Section 5.3. “Description” provides the purpose of the MS. “Resources” lists the REST resources that the service is responsible for. “REST requests” defines the type of REST requests supported for each resource, and how resources are to be represented. “Additional details” provides functional requirements for the service. “Deployment” states deployment details; for instance, that the micro-service will run in a Docker container, the port the MS listens on and the URL used to send this MS requests. An alternative to this format would be to use OpenAPI (Swagger) specifications [24] but we would need to add information necessary to implement the server logic, 3rd party dependencies, etc. The format we provide is more compact.
@@ 3.1 Creating MSBA specifications
To evaluate LLM Agents on MSBAs, we created 8 MSes across 2 MSBAs. We created our own MSBAS for two reasons. First we want to make sure that the LLMs have not been contaminated - have not seen these MSes before. Second, we want to compare LLM’s code generation capabilities on MSBAs of increasing specification complexity, but still small enough to focus on core MSBA challenges and not on extraneous issues. We built the MSBA specifications using the following process:
@@
    When testing and finding an error, we checked the cause of the error. If it was a mistake in the GT implementation, we fixed the GT implementation. If the test case was not faithful to the spec, we corrected the test case. If the spec was ambiguous or incorrect, we fixed the spec.
@@
The specifications and list of unit tests for each MSBA can be found online [^2] https://github.com/LLMs4code/LLMs4MSBA.
@@ 3.2 The complexity of an MSBA specification
We measure the complexity of an MSBA in several ways:
@@
    Resources: the resource complexity. A resource is either a primitive data type (int, string,…), a structure, where each field is a resource, or a collection of resources. If a resource is a single primitive it counts as a single primitive resource. If it is a collection, it counts as a single collection resource. If it is a structure with $k_{1}$ fields that are primitive or structures and $k_{2}$ fields that are collections, it counts as $k_{1}$ primitive and $k_{2}$ collection resources. Structures or containers containing embedded structures or containers are recursively counted on their own. We denote the resource complexity by X-Y where X is the number of primitive resources and Y is the number of collection resources in the application as defined by the spec [^3] Sometimes a resource is computed dynamically - the data is not actually stored. In that case we do count the resource. If a resource does not represent data but is used to take action (e.g.,‘/turn-light-on’) it also is not counted at all. .
  3. 3.
@@
The specification size indicates the number of details the LLM must contend with. The number of dependencies and packages indicates the amount of external knowledge required by the LLM. The number of resources indicates the data management complexity. These metrics are not exact; its possible to have a large specification that is simple to implement and a small specification that is difficult to implement. Similarly not all external dependencies are equal; some express more intricate interfaces than others. Nonetheless, these metrics give a rough estimate on the complexity of an MSBA spec and are easy to compute. The LLM-as-a-Judge serves as a validation for the other metrics, and it also explains why it rated the spec with the given score, providing further insight into the complexity of the spec. Additional metrics could be added, such as the number of conditional requirements/constraints in the specification, similar to measuring Cyclomatic Complexity in programs (see also [15]).
@@ 4 Testing MSBAs
The result of a request to a MS is sensitive to the request history. For instance, the outcome of a GET request is dependent on the preceding POST, PUT and DELETE requests. One needs to test multiple scenarios for a GET to determine whether it functions correctly. Furthermore, the unit tests must also check that the appropriate exceptions are raised when warranted; e.g., when a GET request is issued on an item previously deleted, the status code 404 must be returned whereas if the request provides the wrong media type a status code 415 is returned. Developing adequate unit tests can be time consuming and tedious. To help simplify and make testing of REST APIs more consistent, we developed a REST API testing framework, that consists of 4 modules.
@@
The main module is (i) Unit Test, which is comprised of common unit tests and MS-specific issues the unit tests. It relies on the (ii) HTTP Service Controller to format and invoke the requests according to the HTTP protocol. The common tests are those where the result depends on the pattern of the requests, without concern for their specific input and output values. For instance, if we issue requests POST $X_{1}$, POST $X_{2}$, POST $X_{3}$, DELETE $X_{2}$ and GET, we expect the return value from the GET to be the JSON array containing $X_{1}$ and $X_{3}$ and nothing else. This is true independent of the actual values of $X_{i}$. We rely on this insight to parameterize common unit tests. Parameter values, expected returned values, and any other MS specific values needed to test the MS are defined in the (iii) Database module. The Unit Test module fetches the appropriate values from the DB when invoking a common unit test on a MS; for the example above, it would fetch the values for $X_{1}$, $X_{2}$ and $X_{3}$ relevant for the specific MS.
@@
The common unit tests may still require some customization, and not every such test applies to all MSes. However, the framework greatly simplifies building tests for a new MS. Likewise, we created a reusable parameterized (iv) assertion library that checks that return values satisfy assertions. For example, the assertion assert-fields-contains-value(response: requests.Response,f:str,v) asserts that the response is a JSON array, each element in the array is a struct with field f, the field f is itself a JSON array, and contains the value v.
@@
In addition to common unit tests, we test MS-specific functionality. For example, we check that the Cardholders MS (Figure 1) correctly computes the fine for overdue books. MS-specific tests need to be customized per MS.
@@
Say that an MSBA is made up of 4 MSes: $M_{1}$, $M_{2}$, $M_{3}$, and $M_{4}$. Let $M^{GT}_{i}$ be the GT implementation of $M_{i}$ and $M^{A}_{i}$ be the implementation of $M_{i}$ generated by the LLM agent. To test $M^{A}_{2}$ we run the MSBA using the implementations $M^{GT}_{1}$, $M^{A}_{2}$, $M^{GT}_{3}$, and $M^{GT}_{4}$ and then test the application. In this way we isolate the test failures due only to $M^{A}_{2}$.
@@ 5.1 Methodology
We developed an agentic system built on top of LangGraph [1] and given in Figure 2 to generate and test the microservices (code) of an MSBA given its specification [^4] We refer to each node of the LangGraph as an agent for simplicity of discussion. However, it is commonly accepted to refer to the entire LangGraph as a single agent, and each node of the graph as a function that is part of the agent. . The Init Agent (1) initializes and setups the graph. The Code Agent then (2) generates code for a MS, based upon the specification. Once this code is generated, it passes control to the Docker Test Agent. This agent first (3a) generates a requirements file for the generated code. It then (3b) sets up a Docker container to execute the LLM generated code $M$ and Docker containers to execute all the other MSes using the GT code for each of them. When these containers are up and running the agent (3c) executes the test script on the application and logs any run-time or test errors. Finally it (3d) shuts down all the containers and frees resources before passing control to the Evaluate Agent.
Figure 2: Workflow for generating and testing an MSBA. The blue arrows indicate control flow. The dotted black arrows indicate writing and reading data. Colored boxes denote LLM interactions.
The Evaluate Agent (4a) checks the logs produced by the Docker Test Agent. If there were no errors, then (4b) it records the result and terminates. Otherwise (4c) it checks if the system has reached the maximum number of code-generation trials. If yes, (4b) it records the result and terminates. Otherwise it passes control to the Regen Agent, who (5a) asks the LLM to reflect [27] on the errors and produce a plan to fix the generated code to prevent the top 5 errors. Because of the context window limitations of LLMs, and because the runtime may produce a very long trace of each error, we use a script to remove parts of the error messages that are not useful to debug the issue. For some LLMs, we need to further limit the size of the error messages to fit into the context window. (5b) The original code together with the result of the reflection is then provided to the LLM, which is instructed to generate code that repairs the MS. Control then passes back to the Docker Test Agent (3a) to test the regenerated code. The entire process is repeated for each MS. The prompts used by the agents are given in Appendix A.
In summary, the agentic system automates the entire process from code generation to testing to repair by invoking multiple agents. It uses LangGraph to orchestrate the flow, runs the generated code safely in Docker containers, and uses reflection to improve results.